prettygirl15:

alvarez market is having a sale on fruits and vegetables. the table below shows the price for each item. the sale only lasts this weekend item cost lemons 5 for $4 grapes 4 pounds for $10 bananas 5 for $6 apples 5 for $8 small oranges 8 for $5 how much will it cost for 6 lemons? how much will cost for 6 pounds of grapes? how much will 7 bananas cost? how much will 4 apples cost? how much will 10 smalls oranges cost?

prettygirl15:

Rezo:

ok for the lemon one youll have to divide to find how much one lemon costs so 4/5 = ?

yourdootcome:

banana cost = 6/5 (unit price) * 7 = 8.4 4 apples = 8/5 (unit price) * 4 = 6.4 10 oranges = 5/8 (unit price) * 10 = 6.25

Rezo:

@yourdootcome wrote:
banana cost = 6/5 (unit price) * 7 = 8.4 4 apples = 8/5 (unit price) * 4 = 6.4 10 oranges = 5/8 (unit price) * 10 = 6.25
dont u have to walk them through the problem?

yourdootcome:

@rezo wrote:
@yourdootcome wrote:
banana cost = 6/5 (unit price) * 7 = 8.4 4 apples = 8/5 (unit price) * 4 = 6.4 10 oranges = 5/8 (unit price) * 10 = 6.25
dont u have to walk them through the problem?
ok

Rezo:

i was already tryna do that but im pretty sure theyre afk

yourdootcome:

basically you find the unit price (money over quantity) and you multiply the number of fruits you buy in the problem